2080.1 7

Basic Statistics2080.15 marksNumerical: expected value and varianceExpected value and decision making

A jewelry dealer is interested in purchasing gold necklace for which probabilities are 0.18, 0.22, 0.33 and 0.27 respectively that it will be sold for a profit of Rs. 5000, Rs. 8000, 6000 and sell for a loss of Rs. 3000. Find expected profit and variance of profit.

A jewelry dealer is interested in purchasing gold necklace for which probabilities are 0.18, 0.22, 0.33 and 0.27 respectively that it will be sold for a profit of Rs. 5000, Rs. 8000, 6000 and sell for a loss of Rs. 3000. Find expected profit and variance of profit. [5]

  • Outcomes (profit/loss): Rs. 5000, Rs. 8000, Rs. 6000, and a loss of Rs. 3000 (i.e. $-3000$) - Probabilities: 0.18, 0.22, 0.33, 0.27 Let $X$ = profit in Rs. $X$ 5000 8000 6000 -3000 ------------------------------ $P(X)$ 0.18 0.22 0.33 0.27 Check: $0.18 + 0.22 + 0.33 + 0.27 = 1.00$ ✓ --- $$E(X) =...
Up nextFit a Poisson distribution to the following data. | Defects (X = x) | 0 | 1 | 2 | 3 | 4 |